40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog, Page 2 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  17967   Wed Nov 8 15:14:39 2023 jeffUpdateGeneralAccelerometers

I borrowed a small accelerometer from 40m and walked it to the cryo lab. It is pictured in attachment 2 in its box.

There are two accelerometers I can see on the table in the cryo lab. They are both pictured in attachment 1

Attachment 1: image1(2).jpeg
Attachment 2: image0(3).jpeg
  17966   Tue Nov 7 17:34:36 2023 yutaSummaryGeneralBeam centering strategies around vertex

Here are some beam centering strategies...

Beam height around PRC:
 - From our in-vac inspection today, we noticed that the beam is high on TT2 and PRM, and was slightly clipped (or maybe not) by the earthquake stop of TT2.
 - According to the beam spot position measurements (Attachment #1; includes measurements taken today), beam spot on PRM and PR2 are vertically mis-centered by +1 cm, but vertically pretty centered on PR3, BS and ITMs.
 - These imply we need to play with TT1 and PR2 to level the beam inside the PRC.
 - This is consistent with the height issue we found during the last vent on POP_SM4 (1 inch mirror in front of LO1). We had to place a spacer to make POP_SM4 high to extract the POP beam coming from PR3 to PR2 (40m/16832). So we should be able to remove the spacer after the fix.
 - This fix should be done iteratively without losing flashing in the arms, sacrificing BHD alignment.

Horizontal beam mis-centring on PR3:
 - According to the beam spot position measurements, beam is vertically centered, but horizontally mis-centered by ~1.5 cm at PR3.
 - This is consistent with beam clipping we have in Y green transmission in YAW (Y green is clipped in YAW at PR3).
 - Since the beam is horizontally centered on PRM and PR2, we probably need to physically move PR3 to center the beam and to un-clip Y green transmission.

Beam mis-centering around LO and AS:
 - After these fixes, we probably need to re-align AS path and LO path, as the beams are not centered on LO1, LO2, AS1, AS4.
 - Hopefully these will improve LO-AS mode matching...
 - These probably will not require physical moving of the suspensions.
 - AS RF path after AS2 (90:10 splitter) is also clipped in YAW. This issue also needs to be addressed when re-aligning AS beam.
 - It is probably better to install BHD platform first, before doing LO and AS beam alignment, as BHD platform changes the balancing of ITMY stack, which has AS1 and AS4 suspensions.

Attachment 1: BeamSpotMeasurement_summary.pdf
  17965   Tue Nov 7 16:43:49 2023 yutaSummarySUSUntripping software watchdog

I made a script to untrip software watchdogs.


It is available from the blue "Damp" button in the watchdog screen attached.

Happy in-vac work!

Attachment 1: Screenshot_2023-11-07_16-46-40_WD.png
  17964   Tue Nov 7 16:29:23 2023 RadhikaUpdateGeneralPhotos and e-level measurements of ITMY/BS/ITMX tables

[Yuta, Vittoria, Radhika]

All photo documentation of the open chambers can be found on the 40m google photos account (foteee@gmail.com). Here are the albums for the ITMY, BS, ITMX tables.

We opened each chamber and captured photos of each suspension below. The photos on the google drive have the name of the suspension in the description.

ITMY table: ITMY, AS1, AS4, SRM
BS table: LO2, SR2, BS, PR3, PRM
ITMX table: ITMX, LO1, PR2

The digital tilt-meters (henceforth called e-bubbles) were assigned a label for each table by Yuta. We zeroed their level readings on the PSL table, assuming it is the most level surface around. We placed the e-bubbles as centered as possible on each table keeping distance from the suspensions. *Note: we didn't think to align the e-bubble with the tapped holes of the breadboard, so next time we'll make sure to do this.* These readings can be seen in Attachments 1-3 (ITMY, BS, ITMX).

Each chamber was closed and nothing was moved from any of the tables.

Attachment 1: IMG_5882.JPG
Attachment 2: IMG_5904.JPG
Attachment 3: IMG_5929.JPG
  17963   Mon Nov 6 23:37:13 2023 KojiUpdateGeneralAux Laser: Interlock SWs and fiber

Thanks to Murtaza, now we have two emergency switches for the aux laser. One is on the PSL table, and the other is in the staging HEPA booth.

The fiber was put in a plastic spiral tube for protection and delivered to the staging HEPA booth. I confirmed that the light was coming out from the fiber but haven't checked the output power.

Attachment 1: PXL_20231107_073229088.jpg
Attachment 2: PXL_20231107_073216717.jpg
  17962   Mon Nov 6 21:32:29 2023 PacoUpdateCDSfb1 alarm

[Koji, paco]

First, ssh to fb1 and running sudo dmidecode -t1 we get

# dmidecode 3.2
Getting SMBIOS data from sysfs.
SMBIOS 2.8 present.

Handle 0x0001, DMI type 1, 27 bytes
System Information
    Manufacturer: Supermicro
    Product Name: Super Server
    Version: 0123456789
    Serial Number: 0123456789
    UUID: 00000000-0000-0000-0000-0cc47a699734
    Wake-up Type: Power Switch
    SKU Number: To be filled by O.E.M.
    Family: To be filled by O.E.M.

A quick google search on supermicro superserver "red LED blink" suggests fan failure.
After pulling out fb1 we confirmed the middle of three fans is dead.
The fan is a NIDEC Ultraflo (V80E12BHA5-57), 12 VDC, 0.6 A, 80x80x5 mm (see Attachments #1/#2).

We left the FB1 pulled out, but the dead fan and the lid were left as they were to keep the current airflow condition.

Attachment 1: PXL_20231107_055416153.MP.jpg
Attachment 2: PXL_20231107_060156045.jpg
  17961   Mon Nov 6 21:21:30 2023 KojiUpdateCDSfb1 alarm
  • FB1 has this "i" LED blinking and continuously beeping. What's wrong? 
  • Also, one of the UPS at the bottom of 1X6 has the error "F06".
  • Speaking about UPSs, the control room UPS has a continuous red light and frequent beeping.
Attachment 1: PXL_20231107_051924435.jpg
Attachment 2: PXL_20231107_063145734.jpg
Attachment 3: PXL_20231107_053401989.jpg
  17960   Mon Nov 6 12:27:39 2023 yutaUpdateASCBeam spot position measurements at low power

Beam spot position measurements were done for the first time in air pressure at low power.
Dither amplitudes were increased to overcome lower SNR, but still the uncertainties are large.


# Optic    LSCDoF     freq.(Hz)    ampl. (counts)    gpstime    Opt. gain (counts/nm)    Opt. gain_std    v (mm)    v_std    h (mm)    h_std
ETMY    YARM    211.11    1000    1383331863    64.55    3.78    -4.33    0.37    0.47    0.40
ITMY    YARM    211.11    1000    1383331936    65.15    2.67    -0.97    0.17    -0.20    0.11
ETMX    XARM    211.11    1000    1383331957    59.96    2.10    3.40    0.30    -1.33    0.18
ITMX    XARM    211.11    1000    1383332021    62.73    2.35    0.88    0.15    0.38    0.19
BS    MICH    211.11    3000    1383333747    0.02    0.00    3.21    3.90    -3.33    2.26
PRM    PRY    309.21    3000    1383335425    6.23    6.28    13.55    29.00    -2.05    2.82
PR2    PRY    309.21    3000    1383335499    7.57    4.51    7.70    7.09    3.67    12.17
PR3    PRY    309.21    3000    1383335562    5.88    1.17    3.60    10.33    -16.91    15.18
LO1    HPC    113.13    3000    1383336782    0.05    0.01    9.12    3.44    6.50    5.91
LO2    HPC    113.13    3000    1383336842    0.05    0.01    14.51    6.63    21.34    5.13
AS1    HPC    113.13    3000    1383336901    0.05    0.01    2.62    3.36    14.91    2.70
AS4    HPC    113.13    3000    1383336963    0.05    0.01    -5.84    3.28    14.56    3.47

Locking configurations:
 - YARM and XARM can be locked by increasing the trigger matrix C1:LSC-TRIG_MTRX elements by a factor of 10.
 - MICH can be locked by changing AS55 demodulation phase C1:LSC-AS55_PHASE_R from 2.1 deg to 92.1 deg (why?). Play with trigger matrix C1:LSC-TRIG_MTRX so that it does not trigger filters.
 - PRY can be locked by changing REFL55 demodulation phase C1:LSC-REFL55_PHASE_R from 76.02 deg to 166.02 deg (why?) and by increasing the PRCL gain by a factor of 10, C1:LSC-PRCL_GAIN from -0.04 to 0.4. REFL55 is glitchy and making the measurement uncertainties big. Play with trigger matrix C1:LSC-TRIG_MTRX so that it does not trigger filters.
 - ITMY single bounce-LO can be locked by increasing the gain by a factor of 10, C1:HPC-LO_PHASE_GAIN from 1.5 to 15.
 - These changes are summarized in the following script.


Attachment #1 is the current PRFPMI BHD alignment, and Attachment #1 is the history of the measurements.

Attachment 1: Screenshot_2023-11-06_12-31-32_PRFPMIBHDaligned.png
Attachment 2: BeamSpotMeasurement_summary.pdf
  17959   Sun Nov 5 14:29:19 2023 aaronUpdateEquipment loanN2 regulator returned to cryo

I retrieved the nitrogen pressure regulator from x arm vacuum cabinet and returned it to cryo. Reminder to replace the 40m regulator for the next vent.

  17958   Fri Nov 3 17:16:14 2023 PacoConfigurationIOOLow power IMC locked

[Paco, yuta]

We continued with the alignment recovery by looking for arm cavity flashing. To do this it was very helpful to restore the OpLev centering and use the AUX lasers -- After some further alignment on both arms, we just needed to increase the LSC gains (XARM and YARM) by a factor of ~ 1.5x and lower the TRX/TRY DC matrix elements for triggering from 1.0 to 10.0 in order to engage the IR locks (See Attachment #1).
Attachment #2 shows the screenshot of MEDM screens after both arms are aligned with low power mode.

 - Measure beam spots in low power mode
 - In-vac work to center the beams on vertex optics

Attachment 1: lowpower_arm_lock_20231103.png
Attachment 2: Screenshot_2023-11-03_17-39-31_botharms.png
  17956   Fri Nov 3 17:05:10 2023 PacoConfigurationIOOLow power IMC locked

[Paco, yuta]

We opened the PSL shutter and verified the power behind the shutter is 92 mW. Then, we changed the IMC autolocker parameter file to lower the transmission threshold from 5000 counts to 300 using the instructions from the wiki. We then slightly aligned MC2 in YAW and maximized the flashing in transmission and also changed the parameters for the inputGain slider by +20 dB in both unlocked and locked states so we compensate for the decrease in optical gain. The changes have been pushed here.

After these changes in the config file we ssh into optimus and restart the docker container running the autolocker script in this case by running:

sudo docker restart scripts_AL_MC_1

  17955   Fri Nov 3 13:10:58 2023 KojiUpdateVACFRG1/2/5 recalibrated at 1atm

As we found that some of the FRGs were not well calibrated, Using the driver menu, I've recalibrated the FRGs exposed to 1atm right now.

I expected that the FRGs with the calibration drifted off shows lower or higher pressure like FRG5 which showed 1000 Torr at 1atm.
However, our main FRG1 did not show significant change from 760 Torr even when JC introduced more vent air this morning. This is annoying.

We need to check if the FRG1 is failing, or this is the expected behavior.

Attachment 1: FRG_CAL.png
  17954   Fri Nov 3 10:18:21 2023 JCUpdateVACBHD OMC Vent

This morning, the goal was to open up the chamber doors. Yesterday we attempted, but we did not succeed. Here is a list of the possibilties what we were thinking at the end of yesterday:

  • The annulus pressure was not yet at atmosphere and the door will remain stuck until then.
  • The chamber doors are just sticking to the O-Ring and we need to knock it with a rubber mallet as said in elog 15649.
  • The Main Volume Pressure is NOT at atmosphere and FRG1 AND p1a is showing incorrect pressures.

After speaking on the phone with Jordan, he recommended that we should open VV1 to room air and listen if air is entering the main volume. I did so and air began to rush in.

  17953   Thu Nov 2 14:18:04 2023 yutaSummarySUSNew MEDM screens and misalignment scheme in the era of software watchdogs

New watchdog and suspension summary MEDM screens were made, and IFO_ALIGN screen was updated.
We also implemented new misalign/restore scheme
using C1:SUS-$(OPTIC)_BIASPIT_OFFSET instead of using C1:SUS-$(OPTIC)_TO_COIL OFFSETs in matrix elements.

Script for generating screens with multiple suspensions:
 - A script was made to help generating MEDM screens with multiple optics. A template adl file is used to place each optic in a defined grid.


New misalin/restore scheme:
 - Instead of turning on OFFSETS in C1:SUS-$(OPTIC)_TO_COIL matrices in pitch, C1:SUS-$(OPTIC)_BIASPIT_OFFSET is now used for misaligning.
 - The optic is considered to be misaligned if C1:SUS-$(OPTIC)_PIT_OFFSET and C1:SUS-$(OPTIC)_PIT_BIAS read different value. This can happen with BIAS(PIT|YAW)_OFFSET, C1:SUS-$(OPTIC)_(PIT|YAW)_OFFSET_ON being 0, C1:SUS-$(OPTIC)_BIAS(PIT|YAW)_GAIN being 0 etc.
 - When restoring, these channels are turned on in both pitch and yaw to make sure it is restored.
 - Updated script lives in


New MEDM screens:
 - New C1SUS_WATCHDOGS.adl (Attachment #1), IFO_ALIGN.adl (Attachment #2; does not look new) and C1SUS_SUMMARY.adl (Attachment #3) are attached.
 - In the attached watchdog screen,
   - LO1 is aligned but tripped, not ready to damp because of PD RMS values are larger than MAX_VAR. C1:SUS-LO1_LATCH_OFF=1 and C1:SUS-LO1_LOGIC=1.
   - LO2 is aligned and tripped, but ready to damp. C1:SUS-LO2_LATCH_OFF=1 (this being 1 means it is tripped and not untripped) and C1:SUS-LO2_LOGIC=0 (ready to damp).
   - AS1 is aligned and tripped, but ready to damp. C1:SUS-AS1_LATCH_OFF=0 (this being zero means someone untripped, but does not mean any other) and C1:SUS-AS1_LOGIC=0.
   - AS4 is misaligned and damped. Misaligned meaning (C1:SUS-AS4_PIT_OFFSET != C1:SUS-AS4_PIT_BIAS) || (C1:SUS-AS4_YAW_OFFSET != C1:SUS-AS4_YAW_BIAS).
   - C1:SUS-$(OPTIC)_LATCH_OFF cannot be turned in this screen to show that it does not mean much.
 - For the suspension summary screen, I intentionally kicked AS4 to make PDVARs and alignment PIT offset red. Note that screenshot was taken after the vent, and some oplev values are also red.

Attachment 1: Screenshot_2023-11-02_12-37-55_New_Watchdogs.png
Attachment 2: Screenshot_2023-11-02_14-24-20_IFOALIGN.png
Attachment 3: Screenshot_2023-11-02_14-15-50_SUSSUMMARY.png
  17952   Thu Nov 2 12:06:49 2023 JCUpdateVACBHD OMC Vent

After swappiong out the cylinders, I went to review the pressure to check the rate of the N2, but It looks like we already hit atmospheric pressure?

I'm assuming the room air must have entered the system some how while I was swapping the cylinders. I must've have made a mistake, but I don't understand how room air could've entered while the main volume was isolated during the swap. Could somebody review the steps I took and let me know where I made a mistake? If there is no issue with the procedures, I must have made the mistake and not have noticed. There was a steep rise in the pressure after 200 Torr.


Attachment 1: Vent_11-2-23.pdf
  17951   Thu Nov 2 08:59:11 2023 JCUpdateVACBHD OMC Vent

[Murtaza, JC]

We started preperation of the vent at 7:30 am.

First, to avoid room air entering the system, Murtaza and I purged the line to see if there were any leaks. It turns out that the regulator was the issue! We used soap and water when purging the line to check if we could visibly see an air bubbles, but none were apparent. Yet, there was still a steady hissing somewhere and couldn't figure out where it was coming for. It turned out to be from the control knob of the regulator. I pulled out another regulator from the Vac Cabinet on the X-End, and it worked out perfect.


At 8:30

We finished the leak hunting and moved to begin backfilling the IFO with N2. We are letting in 10 PSI of pressure right now and the Vacuum pressure is steadily increasing approximately 0.02 Torr every 10 min. We are doing great so far

  17950   Wed Nov 1 19:13:48 2023 KojiUpdateSUSSUS medm screen update

We noticed that the alignment offset indicator of the suspension screen only responded to the ON/OFF switches. It was the source of confusion.
What we wanted was for the indicators to react to the switches and gains of the internal filterbanks, in addition to the external ON/OFF switches.

This was done by editing SUS_SINGLE_OVERVIEW.adl in /opt/rtcds/userapps/release/sus/c1/medm/templates/NEW_SUS_SCREENS

The indicator logic was updated like

"dynamic attribute" {

Now the indicators will become green only when the offset switch is ON, the internal input/output switches are both on, and the gain is non-zero.

Attachment 1: Screenshot_2023-11-02_02-13-17.png
  17949   Wed Nov 1 19:10:17 2023 yutaSummarySUSSoftware watchdog implemented in BHD optics, and modified other ones

[Koji, Yuta]

We implemented software watchdog for BHD optics.
In order to make the logic the same for all of the optics, software watchdogs for vertex optics, ETMX and ETMY were also updated.
Now all the OSEM PD VAR are calculated from the fast channels INMON, and turns off suspension damping loops only when tripped.

Software watchdog for BHD optics:
 - Created a template db file for c1sus2 optics to do software watchdog. Now C1:SUS-<OPT>_<OSEM>PD_VAR are calculated based on RMS of C1:SUS-<OPT>_<OSEM>SEN_INMON, and turns off SUSPOS, SUSPIT, SUSYAW, OLPIT and OLYAW dampling loops (but not alignment offsets nor disable coil outputs C1:SUS-<OPT>_<OSEM>_COMM) if triggered.

 - Ran the following for <OPT>=AS1, AS4, LO1, LO2, SR2, PR2, PR3 to create db files for BHD optics.

python /cvs/cds/caltech/target/c1susaux2/createDbFromExcel.py <OPT>

 - Ran the following on c1susaux2

sudo systemctl restart modbusIOC.service

Software watchdog for vertex optics and ETMX:
 - Found that, for ITMX, ITMY, BS, PRM and SRM, oplev loops were not turned off even if they trip. They also disable coil outputs (ETMX as well).
 - Added the following to respective db files.

    field(DESC,"Turn off OPLEV PIT SUS damping")
    field(OUT,"C1:SUS-<OPT>_OL_PIT_SW2S PP NMS")

    field(DESC,"Turn off OPLEV YAW SUS damping")
    field(OUT,"C1:SUS-<OPT>_OL_YAW_SW2S PP NMS")

 - Also commented out five lines below (ETMX as well) so that it does not turn off the coil output switch (we want them to be on so that alignment offsets are not turned off).

    field(DESC,"Fanout for pushing all buttons")
#    field(OUTA,"C1:SUS-<OPT>_UL_COMM PP NMS")
#    field(OUTB,"C1:SUS-<OPT>_LL_COMM PP NMS")
#    field(OUTC,"C1:SUS-<OPT>_UR_COMM PP NMS")
#    field(OUTD,"C1:SUS-<OPT>_LR_COMM PP NMS")
#    field(OUTE,"C1:SUS-<OPT>_SD_COMM PP NMS")

Software watchdog for ETMY:
 - Copied what was done for ETMX to ETMY db file. NOTE that it is /cvs/cds/caltech/target/c1auxey1/ETMYaux.db, and not c1auxey without 1 (40m/17948) !!!!!
 - Ran the following on c1auxey1

sudo systemctl restart modbusIOC.service

 - Make a new watchdog MEDM screen
 - Make a new suspension summary MEDM screen
 - Update the script to untrip watchdogs
 - Update the scprit to align and mis-align suspensions (use C1:SUS-$(OPTIC)_BIASPIT_OFFSET instead of using sneaky offset in coil output matrix?)

  17948   Wed Nov 1 18:52:28 2023 KojiUpdateSUSc1auxey1 confusion

[Koji, Yuta]

Yuta was working on the watchdog unification for all the suspensions. He reached ETMY's one and tried to restart the ETMY's SUS slow machine.

We initially thought that the machine is c1auxey, BUT the correct name is c1auxey1 (horrible name). We lost ~30min because of this confusing name.

There was a folder named c1auxey and c1auxey in /cvs/cds/caltech/target. This was the source of confusion. I've compressed c1auxey into c1auxey.tar.gz. target_archive. We will eventually rename c1auxey1 to c1auxey in order to end this confusion.

Then we found that c1auxey1 could not come back after rebooting. This required the following actions.

... logged into c1auxey

$ df

Filesystem        1K-blocks       Used  Available Use% Mounted on
udev                1992120          0    1992120   0% /dev
tmpfs                402404       6044     396360   2% /run
/dev/sda1         113871044  108170996          0 100% /
tmpfs               2012012          0    2012012   0% /dev/shm
tmpfs                  5120          4       5116   1% /run/lock
tmpfs               2012012          0    2012012   0% /sys/fs/cgroup

modbusIOC was not loaded because the NFS filesystem /cvs/cds was missing.

$ sudo mount /cvs/cds

sudo systemctl daemon-reload
sudo systemctl enable modbusIOC.service
sudo systemctl status modbusIOC.service
sudo systemctl start modbusIOC.service

This made the Y-end epics service back in operation.

  17947   Wed Nov 1 18:10:03 2023 KojiUpdateGeneralReady for the vent

We are ready for the vent


  • I've closed all the shutters for the four lasers.
  • The PSL/End green shutters were also closed.


  • Closed V5 to isolate TP3
  • Stopped TP1 then closed V4 (TP2 isolated)
  • Stopped TP3
  • Stopped TP2 manually (Note that the serial connection broken)
  • Closed TP3/TP2 manual backing valves
  • Stopped AUX RP for TP2/TP3
  • Closed main valve V1 / manual valve to the main volume
  • Closed the annuli volme by VA6
  • Closed VM3 to isolate RGA
Attachment 1: Screenshot_2023-11-01_at_18.06.43.png
  17946   Wed Nov 1 17:02:33 2023 PacoUpdateBHDPSL power lowered

I measured 883 mW of PSL power before the PSL Shutter. I then set the waveplate AXIS 1 to 36.82 deg to make the power 93 mW.

Attachment 1: PXL_20231101_235921597.MP.jpg
  17945   Wed Nov 1 12:28:28 2023 yutaUpdateASCBeam spot position measurements before the vent

Beam spot position measurements were done, oplevs aligned.
Attachment #1 shows IFO alignment with PRFPMI BHD aligned.
Attachment #2 shows oplev values.
Attachment #3 shows the history of beam spot position measurements since Saturday.

Updates on beam spot position measurements:
 - PR2 mass and moment of inertia were updated using the numbers from Solidworks model (40m/17937).
 - Script to measure all of them were prepared. This script still requires human to check the alignment and locking status. /opt/rtcds/caltech/c1/Git/40m/scripts/ASC/measureAllBeamSpotposition.py
 - Script to plot the history: /opt/rtcds/caltech/c1/Git/40m/scripts/ASC/plotBeamSpotPositionMeasurements.py


Attachment 1: Screenshot_2023-11-01_12-26-17_PRFPMIBHDAligned.png
Attachment 2: Screenshot_2023-11-01_12-26-51_OplevsAligned.png
Attachment 3: BeamSpotMeasurement_summary.pdf
  17944   Wed Nov 1 10:58:55 2023 PacoUpdateGeneralBHD OMC Vent

Update Wed Nov 1 10:58:48 2023

All vertex suspensions are now under software watchdog logic. I tested all of them similar to what we did yesterday with ETMX.



  17943   Tue Oct 31 21:09:15 2023 KojiUpdateBHDOMC bond rework

We came back to the OMC 24 hours later and found the bonding work yesterday went well.

We cleaned up the OMC further and then placed it on the BHD platform. Now it's ready for the optical action.

Entrance particle count 2 / Exit particle count 4 for 0.5um

(Bond inspection photos will be posted here by JC)

Attachment 1: PXL_20231031_223632753.jpg
Attachment 2: PXL_20231031_223705466.jpg
  17942   Tue Oct 31 17:42:45 2023 PacoUpdateGeneralBHD OMC Vent

[Koji, Paco]


We just need to replicate what was done in (40m/17941) and test.

  17941   Tue Oct 31 17:27:27 2023 PacoUpdateSUSETMX software watchdog

[Koji, Paco, Yuta]
We implemented the ETMX software watchdog "level 1", meaning the logic disables the SUSPOS/SUSPIT/SUSYAW/OLPIT/OLYAW damping loops whenever the UL/LL/UR/LR/SD PD VAR calculated as soft epics channels in /cvs/cds/caltech/target/c1auxex/ETMXaux.db exceed the PD MAX VAR value stored in C1:SUS-ETMX_PD_MAX_VAR. To do this, we add the following set of record entries to the db file, for example for a single (UL) PD we add:

# Soft Watchdog example from ETMXaux.db
    field(DESC,"ETMX UL Sensor Mean")
        field(SCAN,".2 second")
    field(EGU, "Volts")

        field(DESC,"ETMX UL Se")
        field(SCAN,".1 second")
        field(DTYP,"Raw Soft Channel")
        field(INPA, "C1:SUS-ETMX_ULSEN_INMON")
        field(INPB, "C1:SUS-ETMX_ULPD_MEAN")
        field("CALC", "0.5*ABS(A - B)")

        field(DESC,"ETMX UL Sensor Deviation")
        field(SCAN,".2 second")
        field(INPA,"C1:SUS-ETMX_ULPDMon  NPP")
        field(INPB,"C1:SUS-ETMX_ULPD_MEAN NPP")

and the logic is implemented for all loops by adding entries like:

# Software Watchdog level #1: turns off SUS damping based on OSEM variance data
    field(DESC,"Turn off SUSPOS damping")

# (...) # All other loops

    field(DESC,"Fanout to flip SUS damping switches")

    field(DESC,"Turn off SUS damping")
    field(DTYP,"Soft Channel")
    field(SCAN,".2 second")
    field(DOL, "C1:SUS-ETMX_LOGIC")
      field(OUT,"C1:SUS-ETMX_TRIP_ALL.VAL  PP  NMS")

Next, we ssh into c1auxex and run systemctl restart modbusIOC.service

If you haven't done so or the signals are not available in dataviewer, you can add them explicitly into /cvs/cds/rtcds/caltech/c1/chans/daq/C0EDCU.ini

Finally, you may have to run sudo systemctl restart rts-edc_c1sus from c1sus

Attachment 1: watchdog.png
  17940   Tue Oct 31 17:27:09 2023 KojiUpdateSUSETMX POS unstable when misaligned

[Koji Paco]

We decided to invert the signs of the misalignment offset for ETMX to see it changed the stability of the damping servo when the sus was misaligned.
And actually it did.
The first figure is the individual OSEM outputs while 1) aligned 2) misaligned as before 3) misaligned with the new offsets.
The second figure shows the variations of the OSEM values. We can see that the numbers went up when misaligned and during misaligned state (due to instability). These numbers went down to nominal level quickly when the offset signs were inverted. smiley

Attachment 1: ETMX_misalign1.png
Attachment 2: ETMX_misalign2.png
  17939   Tue Oct 31 16:25:29 2023 KojiUpdateSUSETMX POS unstable when misaligned

I noticed that there was an oscillating spot on the ITMX face camera image. I tracked the source down and found it was the motion of the ETMX.
It turned out that the ETMX pos loop was not stable when it was misaligned. It is very dangerous. We would have the wire broken after many hours of a mirror swinging like this.

This is exactly why we want a functioning watchdog.

Attachment 1: Screenshot_2023-10-31_16-24-54.png
  17938   Mon Oct 30 19:20:16 2023 KojiSummaryGeneralBHD Inventory

Prepared the following components around the staging clean booth:

  • SOS 2"->3" meal sleeve for thick optic (Attachment 1)
    • To mount 1/2" optic we will need two OD 2" (ID 1.8") x 0.125" Al spacers. -> need to be manufactured
  • Other optic mounts (Attachment 2)
    • Vented screws / 1"&2" lens mounts / 1" Polaris mirror mounts
Attachment 1: 46B2767C-E6C1-46D0-B7BD-0475301B241A_1_201_a.jpeg
Attachment 2: 636D6127-B245-42E1-AF05-D4A61D244170.jpeg
  17937   Mon Oct 30 19:11:58 2023 KojiUpdateASCMoIs for the BHD SOSs

From the ITMX window, I confirmed that PR2 is using the original metal sleeve and the optic is 3/8" thick. (See Attachment 1)

Based on this, I made a SW model to check the mass and the IoM.

== 0.375 inch optic / original sleeve ==

- Mass = 201 grams
- Moment of Inertia (grams mm^2)
Lxx = 1.01e5 (Pitch)
Lyy = 9.41e4 (Yaw)
Lzz = 1.79e5 (Roll)

Attachment 1: PXL_20231031_004416449.NIGHT.jpg
  17936   Mon Oct 30 18:04:46 2023 PacoUpdateSUSBS Oplev check

I measured the Oplev PIT/YAW OLTFs and adjusted the gains to get 0.1 to ~ 4 Hz feedback as is recommended in Gautam's thesis (Fig 3.3 and discussion thereafter). Attachments #1-2 show the measured OLTF and simple expected model using the foton filters and plant. Attachment#3 shows the final filter modules and gains after the changes were done.

Next steps:

  • Repeat for PRM, ITMs and ETMs
Attachment 1: BSOL_OLTF_measuredScreenshot_2023-10-30_18-11-12.png
Attachment 2: IDEAL_BSOL_OLTFScreenshot_2023-10-30_18-10-19.png
Attachment 3: FINALBSOL_screenScreenshot_2023-10-30_18-10-52.png
  17935   Mon Oct 30 15:47:16 2023 KojiUpdateBHDOMC bond rework

[JC Koji]

We successfully applied EP30-2 on the OMC #1.

  • We brought a scale with 0.01g resolution from Yuta's desk.
  • The Black and Decker toaster oven was brought from the bake lab (it was already returned)
  • Made two AL cups for epoxy mixing
  • Brought an SS spatula from the bake lab.
  • Checked which bracket has the 50% delamination by lifting the OMC --> it was the one close to the QPDs (Attachment 1)
  • Used one of the EP30-2 tubes with an expiration date of Nov 1, 2023 (in 2 days!) (Attachment 6).
  • Disposed one push from the glue applicator tube into one of the Al cups. Then, poured 7g into the other Al cup.
  • Added 0.35g of silica bead powder.
  • Mixed -> painted a few drops on a shell of Al foil -> Bake in the oven at 200F for 15min.
  • JC and I confirmed the baked test piece was nicely cured. No stickiness. Very crisp.
  • Applied the bond on the side of the 50% delaminated bracket. (Attachments 2/3).
  • I could confirm that the glue was sucked into the gap somewhat. The action was still going. (Attachment 4).
  • Now, we proceeded to the cable bracket. JC applied the layer of glue to the two brackets.
  • Then, the cable bracket was placed -> We checked how the upper part was wet. It was very nicely done! (Attachment 5)
  • We've placed some weight on the cable bracket so it does not move. In reality, the bracket was well constrained by the four screws and had almost no play to slide. (Attachment 7)
  • Decided to resume tomorrow at 2:30PM
  • The exist particle counting: 0 count for 0.5um particle.
Attachment 1: F71528AC-6985-411F-8463-D5C0CB56D033.jpg
Attachment 2: pxl_20231030_215836571.jpg
Attachment 3: pxl_20231030_215844819.jpg
Attachment 4: pxl_20231030_220137914.jpg
Attachment 5: 9733B2DC-E88D-4B54-AE8F-108196C568C6.jpg
Attachment 6: pxl_20231030_221427003.mp.jpg
Attachment 7: pxl_20231030_220504732.jpg
Attachment 8: pxl_20231030_221410803.jpg
  17934   Mon Oct 30 14:29:20 2023 yutaUpdateASCBeam spot position measurements with updated moment of inertia

Beam spot position measurements were done again using updated estimate of moment of inertia.
Also, LPF for demodulation is applied with sosfiltfilt instead of filtfilt.


# Optic    LSCDoF     freq.(Hz)    ampl. (counts)    gpstime    Opt. gain (counts/nm)    Opt. gain_std    v (mm)    v_std    h (mm)    h_std
ETMY    YARM    211.11    100    1382731658    3601.14    50.48    0.22    0.04    -0.14    0.02
ITMY    YARM    211.11    100    1382731730    3616.83    45.92    0.10    0.02    0.06    0.03
ETMX    XARM    211.11    100    1382731794    1397.71    17.22    0.52    0.10    0.42    0.09
ITMX    XARM    211.11    100    1382731852    1600.21    17.47    1.18    0.20    0.09    0.03
BS    MICH    211.11    3000    1382732022    1.50    0.03    2.61    1.29    -2.60    0.95
PRM    PRY    211.11    300    1382732337    237.32    11.38    8.07    1.06    -2.95    0.30
PR2    PRY    211.11    300    1382732455    354.85    6.44    6.79    0.66    0.62    0.05
PR3    PRY    211.11    300    1382732578    276.86    10.08    0.98    0.07    -16.29    1.17
LO1    HPC    113.13    1000    1382733109    3.70    0.08    7.80    1.32    5.18    0.72
LO2    HPC    113.13    1000    1382733180    3.55    0.53    -14.40    2.99    21.47    5.69
AS1    HPC    113.13    1000    1382733248    3.73    0.18    1.69    1.87    15.67    0.92
AS4    HPC    113.13    1000    1382733325    3.70    0.27    -5.84    1.17    13.26    1.79
SR2    HPC    113.13    1000    1382733385    3.41    0.77    1.46    3.05    6.98    3.75

Attached is the comparison between measurements done on Friday and today.

What we did:
 - Updated the code to use scipy.signal.sosfiltfilt instead of scipy.signal.filtfilt.

Wn=0.0001    # Butterworth filter cutoff will be fs*Wn = 1.6834 Hz for fs=2**14 Hz
#[bbb, aaa] = sg.butter(4, Wn)
sosLPF = sg.butter(4, Wn, output='sos')
#I = 2*sg.filtfilt(bbb, aaa, data*sin)
#Q = 2*sg.filtfilt(bbb, aaa, data*cos)
I = 2*sg.sosfiltfilt(sosLPF, data*sin)
Q = 2*sg.sosfiltfilt(sosLPF, data*cos)

 - Updated the code to use Solidworks estimated mass and moment of inertia in 40m/17930. PR2 mass and moment of inertia was estimated by removing 0.25 inch thick mirror from "0.25 inch optic / original sleeve" model and adding 3/8 inch optic.

 Mass = # in kg
 {'BS': 0.255,
 'ITMX': 0.255,
 'ITMY': 0.255,
 'PRM': 0.255,
 'ETMX': 0.255,
 'ETMY': 0.255,
 'MC2': 0.255,
 'PR2': 0.199,
 'PR3': 0.235,
 'LO1': 0.185,
 'LO2': 0.235,
 'AS1': 0.185,
 'AS4': 0.185,
 'SR2': 0.235}

MoI = # in kg*m**2
{'BS': {'PIT': 0.000106, 'YAW': 0.000106},
 'ITMX': {'PIT': 0.000106, 'YAW': 0.000106},
 'ITMY': {'PIT': 0.000106, 'YAW': 0.000106},
 'PRM': {'PIT': 0.000106, 'YAW': 0.000106},
 'ETMX': {'PIT': 0.000106, 'YAW': 0.000106},
 'ETMY': {'PIT': 0.000106, 'YAW': 0.000106},
 'MC2': {'PIT': 0.000106, 'YAW': 0.000106},
 'PR2': {'PIT': 0.00010091055503756133, 'YAW': 9.391055503756132e-05},
 'PR3': {'PIT': 9.85e-05, 'YAW': 9.999999999999999e-05},
 'LO1': {'PIT': 9.84e-05, 'YAW': 9.14e-05},
 'LO2': {'PIT': 9.85e-05, 'YAW': 9.999999999999999e-05},
 'AS1': {'PIT': 9.84e-05, 'YAW': 9.14e-05},
 'AS4': {'PIT': 9.84e-05, 'YAW': 9.14e-05},
 'SR2': {'PIT': 9.85e-05, 'YAW': 9.999999999999999e-05}}


 - Automate the lock part so that we can run often and do health check of IFO.
 - Check clipping by measuring the power change at each port while dithering.

Attachment 1: BeamSpotMeasurement_summary.pdf
  17933   Mon Oct 30 13:28:28 2023 KojiUpdateBHDEpoxy removal

[JC Koji]

In the preparation of the OMC bond rework, we removed the remaining bond layer on the glass brackets.

Tools: Newly opened acetone, small glass bottle, syringes, Cotton Q-tips (wooden stick), cleaned razor blades


  • The epoxy layer stayed on the glass bracket surface.
  • We first soaked the top of the brackets with acetone, hoping the bond would start to melt.
  • We didn't see much difference after ~5 min.
  • We started scrubbing the layer with the cotton Q-tips. We hardly saw any change.
  • After struggling for ~30 min, some delaminated areas started to peel off, particularly when acetone liquid was applied. (Attachment 1)
  • Kept applying acetone while the surface was scratched with a razor blade. We try not to stress the middle part of the glass bracket as it's the most fragile part.
  • This process had the risk of being messy. Sheets of the lens cleaning paper were laid out around the cable bracket (Attachment 2).
  • All the glue on the brackets was removed. This turned the matte surface shiny and smooth.
  • Removed all the messy paper from the OMC. The fragments of the epoxy layers were removed by Kapton tape, acetone flow from syringes, and IPA wiping.
  • The result looked like Attachment 3.


  • Apply EP30-2
  • The cable bracket seems to keep its position with no position constraint. Nevertheless, we'll place some metal weights around and on the cable bracket.
  • The bracket close to the QPDs has ~50% delamination. The bond is going to be applied there also.
Attachment 1: 5312C29E-22F9-4B94-8B64-31D383A59F0C.jpg
Attachment 2: EED91069-2F51-4455-AD8D-1E62B66FF4D7.jpg
Attachment 3: PXL_20231030_192818520_15.52.00.jpg
  17932   Sun Oct 29 21:09:23 2023 KojiUpdateBHDOMC cable bracket delamination

[JC, Koji]

While adjusting the kinematic mount, we noticed that the OMC cable bracket was sagging from the OMC breadboard. This happened with no specific impact given.

This was the delamination of the epoxy (EP30-2) between the glass brackets and the rectangular metal shims. (Attachments)
There was no real damage to the components.

Also, a small delamination area was found on one of the mass mounting brackets (not pictured).

# This OMC was the first manufactured OMC in 2013, installed at LLO, and taken out last year. This is not the unit to be (possibly) used at Hanford.

The EP30-2 of this era had no established procedure for proper mixing yet. We've experienced frequent delamination in the OMCs and suspension assemblies.

Now that procedures for testing epoxy mixtures in a toaster oven have been established, the frequency of such delamination has decreased.

My assessment for the repair procedure:
- Remove the remaining epoxy layer on the glass side by scrubbing with a cotton swab using acetone.
- Apply EP30-2 on the cable bracket.
- Inject EP30-2 mixture to the mass mounting bracket via capillary action.

Strictly speaking, this OMC is a loan from aLIGO, so I will contact the aLIGO team to see if they have no objections to the repair before we start the work.
In the meantime, we can prepare optics for locking the OMC.

Ed (Oct 30, 2023 8AM): The rework was approved by GariLynn and Gabriele.

Attachment 1: PXL_20231027_182848934.jpg
Attachment 2: 296D06E4-B935-4462-84E4-7FE1CCDA8204_1_105_c.jpeg
  17931   Sun Oct 29 21:01:18 2023 KojiUpdateBHDIntegration of the BHD platform and an OMC

[JC, Koji]

Integration of the BHD platform and an OMC

- Working environment: The BHD platform was placed on the optical table, and the parts and tools were placed on a wagon. (Attachment 1)

- The OMC used: OMC #1 (used OMC taken out from LLO)

- Cleanliness: HEPA booth / Particle counter borrowed from the PSL table. 0~20 count per cubic foot for 0.5um.

- OMC mounting scheme: The kinematic mount with three balls and three grooves is supposed to fix the position of the OMC once it is aligned. 

- OMC mounting procedure: Firstly, the upper parts were fixed to the OMC  (Attachment 2), while the lower parts were placed on the BHD Platform without fixing  (Attachment 3). Then, the OMC was placed on top. By jiggling the bottom parts, the grooves were aligned to the given ball positions (Attachment 4). Because the lower parts have grooves, they still can move before they are fixed keeping the tangent of the grooves and the balls. After the lower parts were aligned with the fixing thread holes well within the through holes, the hexagonal bolts were fastened with a spanner.

Using this OMC, the lower mounting parts for the second OMC slot were also aligned. They will need to be adjusted with the other OMC given later.

It is important to note that when raising or lowering the OMC from the BHD Platform, the fingers should be put on the attachment for the kinematic mounts from the bottom rather than pulling the glass breadboard up by putting the fingers on the glass breadboard.

This is to avoid extension stress on the glass mounting brackets, which can lead to epoxy delamination and/or cracking of the glass parts. Lifting the attachment for the kinematic mounts from below is safe because only compressive forces will be applied to the bonding on the breadboard surface.


Attachment 1: PXL_20231027_172119222.jpg
Attachment 2: A7893E0A-686B-4C52-8983-37E2525A921F.jpg
Attachment 3: PXL_20231027_174506239.jpg
Attachment 4: PXL_20231027_182842030.jpg
  17930   Sat Oct 28 18:07:37 2023 KojiUpdateASCMoIs for the BHD SOSs

I checked Solidworks models to see the mass/moment of inertia for the BHD SOS masses.
The mass/MoI calculation includes 2" glass optic. MoIs were given at the center of mass.

== 0.25 inch optic / original sleeve ==

- Mass = 185 grams
- Moment of Inertia (grams mm^2)
Lxx = 9.84e4 (Pitch)
Lyy = 9.14e4 (Yaw)
Lzz = 1.74e5 (Roll)

== 0.5 inch optic / thick optic sleeve ==

- Mass = 210 grams
- Moment of Inertia (grams mm^2)
Lxx = 9.37e4 (Pitch)
Lyy = 9.52e4 (Yaw)
Lzz = 1.73e5 (Roll)

== 0.75 inch optic / thick optic sleeve ==

- Mass = 235 grams
- Moment of Inertia (grams mm^2)
Lxx = 9.85e4 (Pitch)
Lyy = 1.00e5 (Yaw)
Lzz = 1.80e5 (Roll)

== Ref: 3" x 1" glass optic ==

- Mass = 255 grams
- Moment of Inertia (grams mm^2)
Lxx = 1.06e5 (Pitch)
Lyy = 1.06e5 (Yaw)
Lzz = 1.85e5 (Roll)

I could not find the model for the 3/8" thick optic. The original metal sleeve was designed for 1/4" optic. It is probably possible to put the 3/8" optic into it (with significant misalignment).
It seems that PR2 was installed on Jan 27, 2022. However, both the photograph record and elogs were not detailed enough to identify which metal sleeve was used for PR2.
We need to look into the viewport to identify it.

To fit the 3/8" optic into the thick optic (3/4") sleeve, 3/16" thick ring spacers were required. I have no record of making such sleeves.

The new PR2 optic has a thickness of 1/2". The Solidworks model has the version with 1/2" optic, which was fit into the thick optic sleeve with two 1/8" spacers.
We made a total of 4 thick sleeves. Even if the current PR2 has the original metal sleeve, we are supposed to have one more of the thick sleeves. We still have to find/manufacture Qty 2 of the 1/8" spacers.

  17929   Sat Oct 28 18:04:43 2023 KojiUpdateGeneralEP30-2 kit brought from Downs

EP30-2 kit brought from Downs on Oct 27, 2023. Stephen said he would take care of the property (location) record.

Also, Stephen mentioned that he left the gluing kit in the clean room next to the bake lab. I went there and found the box. It only has two bottles of silica bead powder.

Attachment 1: PXL_20231027_233829648.jpg
Attachment 2: PXL_20231027_233842362.jpg
  17928   Sat Oct 28 02:33:22 2023 yutaUpdateASCBeam spot position measurements with updated demodulation method

Beam spot position measurements were done again using updated demodulation method.
The measurements now also includes statistical uncertainty (Note that uncertanties below do not include uncertainties in actuation, moment of inertia, and coil balancing).


# Optic    LSCDoF     freq.(Hz)    ampl. (counts)    gpstime    Opt. gain (counts/nm)    Opt. gain_std    v (mm)    v_std    h (mm)    h_std
ETMY    YARM    211.11    100    1382518790    3764.95    176.42    -0.23    0.11    0.61    0.05
ITMY    YARM    211.11    100    1382518963    3555.57    89.67    -0.13    0.04    -0.26    0.03
ETMX    XARM    211.11    100    1382519267    1295.22    17.97    -0.35    0.20    -0.98    0.18
ITMX    XARM    211.11    100    1382519351    1456.79    36.07    0.24    0.03    0.47    0.07
BS      MICH    211.11    3000    1382519739    1.60    0.10    1.81    0.68    -2.20    0.40
PRM     PRY    309.21    300    1382520366    256.02    12.97    -7.36    1.21    2.74    1.13
PR2     PRY    309.21    300    1382520438    366.31    24.85    -4.84    0.86    -0.33    0.11
PR3     PRY    309.21    300    1382520585    314.60    13.66    0.62    0.96    -13.57    2.67
LO1     HPC    113.13    1000    1382520845    3.78    0.08    5.19    0.20    3.92    0.11
LO2     HPC    113.13    1000    1382520919    3.76    0.12    -12.98    0.99    18.44    0.70
AS1     HPC    113.13    1000    1382520990    4.07    0.08    0.49    0.12    10.84    0.38
AS4     HPC    113.13    1000    1382521061    4.13    0.10    -3.65    0.14    9.28    0.36
SR2     HPC    113.13    1000    1382521135    4.04    0.07    0.07    0.12    5.30    0.51

See attachment for example spectra during the dither for BS.

What we did:
 - Instead of


for demodulating, the following code was used.

Wn=0.0001    # Butterworth filter cutoff will be fs*Wn = 1.6834 Hz for fs=2**14 Hz
[bbb, aaa] = sg.butter(4, Wn)
I = 2*sg.filtfilt(bbb, aaa, data*sin)
Q = 2*sg.filtfilt(bbb, aaa, data*cos)
mask=int(1/Wn)    # remove first part during LPF time constant in averaging

 - 4th order Butterworth LPF filter with cutoff of 1.6834 Hz was used, and the averaging was done afterwards. Uncertainty was estimated using the standard deviation of LPF-ed data. The first part of data was removed for estimating the average and uncertainty, to wait for the LPF effect to settle down.


 - Get moment of intertia for BHD compound optics from CAD.
 - Automate the lock part so that we can run often
 - Check clipping by measuring the power change at each port while dithering.

Attachment 1: BeamSpotMeasurement_BS_MICH_1382519739.pdf
  17927   Fri Oct 27 17:37:26 2023 yutaUpdateASCBeam spot position measurements with sign

Beam spot position measurements were done again with the correct sign on the mis-centering.
The sign tells you the sign of A2L when you dither in angle, with respect to the sign of your length dither. It does not tell you if the beam spot is on left/right or upper/lower, but it keeps track of which direction the beam is mis-centered.


# Optic LSCDoF  freq. (Hz)   ampl. (counts) gpstime Opt. gain (counts/m) v (mm)   h (mm)
ETMY    YARM    211.11    100    1382471007    3.58e+12    0.17    0.12
ITMY    YARM    211.11    100    1382471069    3.48e+12    -0.02    0.06
ETMX    XARM    211.11    100    1382471097    1.27e+12    -0.60    0.15
ITMX    XARM    211.11    100    1382471162    1.44e+12    0.36    0.66
BS      MICH    211.11    3000    1382471289    1.61e+09    2.00    -2.11
ITMY    MICH    211.11    3000    1382471345    1.58e+09    -0.40    -0.88
ITMX    MICH    211.11    3000    1382471405    1.58e+09    0.68    0.57
PRM    PRY    309.21    300    1382487717    2.94e+11    -7.00    2.06
PR2    PRY    309.21    300    1382487787    4.41e+11    -4.41    -0.41
PR3    PRY    309.21    300    1382487847    3.41e+11    -0.70    12.98
ITMY   PRY    309.21    300    1382487963    3.39e+11    0.10    -0.28
LO1    HPC    113.13    1000    1382488200    3.54e+09    5.05    3.93
LO2    HPC    113.13    1000    1382488279    3.64e+09    -12.48    16.24
AS1    HPC    113.13    1000    1382488340    3.77e+09    -0.72    -10.92
AS4    HPC    113.13    1000    1382488397    3.80e+09    -3.39    9.30
SR2    HPC    113.13    1000    1382488482    3.81e+09    -0.18    -5.77

See attachment for example spectra during the dither for LO1.

What we did:
 - We repeated the measurements in 40m/17925 again, but with the correct sign on the mis-centering.
 - The sign on horizontal and vertical mis-centering was defined by the phase difference in the error signal between position dither and pitch/yaw dither. The phase from the position dither was used as a reference to check the sign of the error signal.
 - In the previous version of the code, demodulation was done using a sin wave generated in the code, but now the script also demodulates the excitation signal, C1:SUS-(optic)_(LSC|ASCPIT|ASCYAW)_EXC, to correctly measure the phase relative to the excitation.
 - The script also now saves the optical gain in BeamSpotPositionMeasurements.txt to keep track of things. Note that, even if the optical gain calibration was off, for example from no notching, it doesn't ruin the beam spot measurements, because we are just measuring the relative length noise difference between the length dither and angular dither.
 - The script now handles different thickness of mirrors (but it does not correctly take into account of compound mirror we have for BHD optics...)
 - The script now do not import math * (I'm just used to pi instead of np.pi).


Both YARM and XARM ASS work.yes

 - Fix demod part of the script to use a low pass filter (I tried scipy.signal.filtfilt, but failed. Probably because the cut off frequency of scipy.signal.butter I used being too low compared with the sampling frequency or something).
 - Get moment of intertia for BHD compound optics from CAD.
 - Check clipping by measuring the power change at each port while dithering.
 - Add uncertainty in the measurements.

Attachment 1: BeamSpotMeasurement_LO1_HPC_1382488200.pdf
  17926   Fri Oct 27 13:30:16 2023 ranaUpdateASCBeam spot position measurements

I looked at this code and it is a little untrustworthy. The demod is not being done correctly so there can easily be some aliasing going on.

in this lab we really, really should never use a moving average instead of low pass filtering.

And why import math instead of using numpy??

  17925   Thu Oct 26 21:37:08 2023 yutaUpdateASCBeam spot position measurements

[Begüm, Paco, Yuta]

Beamspot position measurements were done using A2L, assuming all the coils are balanced.
Beam centering on PR3, LO2, AS1, AS4 are off by more than 1 cm.
ITMX centering (which we ignore in XARM ASS) is pretty good.


# Optic LSCDoF  freq(Hz)   amplitude   gpstime v (mm)   h (mm)
ETMY    YARM    211.11    100.0    1382413915    0.20    -0.20
ITMY    YARM    211.11    100.0    1382413976    -0.01    0.07
ETMX    XARM    211.11    100.0    1382414046    0.08    -0.17
ITMX    XARM    211.11    100.0    1382414280    0.43    0.42
BS      MICH    211.11    1000.0    1382414504    2.00    -2.30
PRM     PRY     309.21    300.0    1382415273    -7.70    -1.85
PR2     PRY     309.21    300.0    1382415349    5.37    0.25
PR3     PRY     309.21    300.0    1382415473    0.85    13.59
ITMY    PRY     309.21    300.0    1382415569    -0.25    0.15
LO1     HPC     113.13    1000.0    1382416132    -5.78    -4.43
LO2     HPC     113.13    1000.0    1382416189    -13.79    18.33
AS1     HPC     113.13    1000.0    1382416253    0.85    12.61
AS4     HPC     113.13    1000.0    1382416315    -4.21    10.65
SR2     HPC     113.13    1000.0    1382416389    -0.38    5.45

 - Lock some interferometer configuration which involves the optic you want to measure (e.g. PRY for PRM).
 - Put a notch filter at the frequency you want to dither in the LSC loop.
 - Dither C1:SUS-(optic)_LSC_EXC and demodulate the error signal to get the optical gain of the error signal.
 - Dither C1:SUS-(optic)_ASC(PIT|YAW)_EXC and demodulate the error signal. Calibrate the demodulated error signal into meters using the optical gain derived above. Calibrate the optic motion using angular actuation efficiency estimated using the method described below. By dividing the length change by angular motion, you get the mis-centering.
 - This method asumes that there is no clipping, all the coils are balanced, POS/PIT/YAW are purely actuated, and optic center is the same as actuation node.
 - For measuring the beam spot positions on ETMs and ITMs, signle arm locking configurations were used.
 - For measuring the beam spot position on BS, MICH configuration was used.
 - For measuring the beam spot position on PRM, PR2 and PR3, PRY configuration was used.
 - For measuring the beam spot position on LO1, LO2, SR2, AS1, and AS4, ITMY single bounce vs LO configuration was used.
 - Attachment #1 shows the spectra of error signals during dither, showing SNR is pretty good.


Estimating angular actuation efficiency:
 - Angular actuation efficiencies were estimated using the following

A_ang = A_lsc * dd * mass / (2*I)

where A_lsc is the length actuator efficiency measured in 40m/17886 and 40m/17918, and

rr = 3*inch/2    # radius of the optic
tt = 1*inch  # thickness (ASSUMPTION!!!; they are random for BHD optics (;x;) (╯°□°)╯︵ ┻━┻)
dd = 1.945*inch # distance between magnets https://dcc.ligo.org/LIGO-D960002
rho = 2.201*g/(cm**3) # density of fused silica (note that BHD optics have 2inch -> 3inch sleeve made of aluminum)
mass = pi*rr**2*tt*rho # mirror mass
I = mass*rr**2/4+mass*tt**2/12 # moment of inertia

 - In summary, they are the following. Note that A_ang is not the physical angle you move, but is what you see by LSC through A2L.

Susp. | LSC meas. nm/count | Gain adj. | AoI deg | POS nm/count | ANG urad/count
BS   +69.54 +2.325 45 +21.15 +4.12
ITMX +14.73 +0.653  0 +22.55 +0.87
ITMY +14.50 +0.659  0 +21.99 +0.86
ETMX +12.20 +0.414  0 +29.47 +0.72
ETMY +10.66 +0.480  0 +22.21 +0.63
MC2  +2.27 +0.105  0 +21.62 +0.13
PRM  -19.00 +0.773  0 -24.58 -1.13
PR2  -36.19 +1.000  0 -18.09 -2.15
PR3  -29.57 +1.000 45 -20.91 -1.75
LO1  +28.25 +1.000  0 +28.25 +1.67
LO2  +11.19 +1.000 45 +15.83 +0.66
AS1  -25.92 +1.000  0 -25.92 -1.54
AS4  +25.82 +1.000  0 +25.82 +1.53
SR2  -16.63 +1.000 45 -23.52 -0.99

 - Check the sign of mis-centering. It is kind of random now.
 - Put uncertainty in these measurements.
 - Include HPC in the LSC screens

Attachment 1: BeamSpotMeasurements_20231026.pdf
BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf BeamSpotMeasurements_20231026.pdf
  17924   Thu Oct 26 19:40:57 2023 RadhikaUpdateASSXARM ASS Working

XARM ASS Fixed (hopefully)

The winning approach was considering T and L loops separately, and adjusting the gain hierarchy. After chatting with Rana, we reconfirmed that the centering length (L) loop should feed back to the cavity optics, and the pointing transmission (T) loop should feed back to the BS. We discussed dithering the BS to generate a pointing error signal, but it turned out this wasn't necessary since a solution was found with just dithering the ITM and ETM. I decided to make the T loop fast and the L loop slow, as was done previously by Anchal.

Attachment 1 shows the final servo gains and output matrix, along with the Striptool showing maximized transmission and suppression of error signals.

T loop

The intuition was to use the ITM T signals (transmission demodulated at ITM dither freq) as a proxy for the BS pointing error, as was done previously:


Next, ETM T signals were fed back to the ETM to maximize transmission. This has always worked:


The signs were chosen based on what suppressed the error signals.

L loop

On paper I worked through how ITM/ETM misalignment shifts the beam spot on both the ITM and ETM. This was mainly a helpful exercise to gain intuition for the centering. I made small angle assumptions and ended up with:

\begin{pmatrix} d_{ETM} \\ d_{ITM} \end{pmatrix} = \begin{pmatrix} r & r \\ r & r-l \end{pmatrix} \begin{pmatrix} \theta_{ETM} \\ \theta_{ITM} \end{pmatrix}

Here r is the radius of curvature of the ETM (57.37cm); l is the length of the arm cavity (40m); d is the displacement from center of the optic; θ is the angular misalignment of each optic.

In practice, we do not care about the centering of the beam on the ITM. So in reality the useful takeaway was that ETM and ITM angular misalignment both shift the beamspot on the ETM by roughly the same amount.



Again the signs were chosen to suppress the ETM L error signals without the T loop on.

Lastly, I chose the T loop gain 0.1 and L loop gain 0.2.

Instructions for running

The servo gains and output matrix have been updated in ASS_Settings_XARM.snap. Just hit "Run ASS".

Attachment 1: XASS_working_TL_loops_2023-10-26.png
  17923   Thu Oct 26 18:34:58 2023 KojiSummaryGeneralOMC #1 / DCPDs / Glass Breadboard are in

[Begüm, Koji]

We brought the OMC #1 / DCPDs / a glass breadboard from the OMC lab to the 40m. They are placed on the wagon next to the staging table.

Attachment 1: PXL_20231027_011853566.MP.jpg
  17922   Thu Oct 26 13:54:08 2023 KojiSummaryGeneralTowards vent on Nov 1st.

= BHD Platform assembly on the staging table =

  • [Done] The assembly will be (mostly) done by the end of Thu (JC)
  • [Done] OMC is going to be mounted on Fri (KA/JC)
  • After the vent, we will bring out the BHD optic from the ITMY chamber.
  • We'll continue to work on the BHD platform alignment on the staging table
    with the optics/optic mounts taken out from the chamber.
    + the OFI components (incl. the machines posts)
  • Asking Dean for the mighty mouse connectors for the devices
    • -> Asked. We need the connectors and pins. Dean is working on the procurement.

= OMC mounting / locking =

  • [Done] Bring the OMC#1 from the OMC lab (KA, Thu afternoon)
  • [Done] Attach the kinematic mounting brackets on the OMC. 
  • Blank OMC breadboard for counterweighting (KA, bring from the OMC lab) -> Brought
  • Locking electronics 
    • [Done] Paco prepared:
      Moku, PDA10 (150MHz), HV Amp for Laser or OMC PZT

    • Prepare locking optics (steering mirrors, post, etc)
  • OMC DCPD prep
    • [Done] Bring the DCPDs (Excelitas C30655) from the OMC lab
    • Remove the cap of the diodes (Asked Dean about the tool)
    • DCPD installation on the OMCs

= IFO work until the vent =

  • [Done] X arm ASS repair (Radhika)
  • FPMI / PRMI locking
    • BS is moving a lot -> fix

= Vent Prep =

  • [Done] Watchdog implementation on all the optics
  • Clipping investigation

= Vent =

  • Table tilt check (digital tilt meter)
  • Open ITMY:
    • pick up components for BHD platform
    • Turn SRM
    • BHD platform installation
    • Counter weight adjustment
  • Open ITMX:
    • PR2 sus optic swap -> optic thickness 3/8" -> 1/2", needs adjustment
  • Open BS / INJ
    • Clipping check
  • Vacuum maintenance
    • Remove TP1 for maintenance
    • TP2 communication error problem
  17921   Thu Oct 26 13:50:14 2023 KojiUpdateGeneralPicomotor / HWP rotator wiring

Wiring configuration for the picomotors and the HWP rotator.
The mighty mouse connectors on the devices are supposed to mate with the cable D2300210.
I'm asking Dean to help us for the mighty mouse connectors.

Attachment 1: 40mBHD_OMC_mighty_mouse_wiring.pdf
  17920   Wed Oct 25 14:03:17 2023 RadhikaUpdateLSCFPMI CARM with CM board

I drew a block diagram to work through the high-bandwidth CARM feedback using the CM board. The goal was to obtain a derivation for the loop algebra used in CM board YARM locking here, applied to CARM locking. Attachment 1 is the diagram with the following blocks:

CIMC: IMC cavity TF (incl. pole, optical gain)
CCARM: CARM cavity TF (incl. pole, optical gain)
FIMC: MC servo board TF
FCMB: CM servo board TF
FLSC: digital CARM controller TF

Below are the independent, decoupled OLTF expressions for each loop (other loops off):

1. GIMC is the OLTF of the IMC-->PSL loop (middle loop in diagram):

G_{IMC} = F_{IMC}\,C_{IMC}

2. GLSC is the OLTF of the slow LSC loop (bottom loop in diagram):

G_{LSC} = G_{IMC}\,C_{CARM}\,F_{LSC}

3. GCMB is the OLTF of the fast CM board loop (top loop in diagram):

G_{CMB} = G_{IMC}\,C_{CARM}\,F_{CMB}

Now that the independent loop OLTFs are defined, we can interpret the loops that have been measured:

1. The LSC slow CARM loop is always measured with the IMC locked, so the measured TF will include the IMC loop supression:

G_{LSC}^{IMC_{on}, CMB_{off}} = \frac{G_{LSC}}{1-G_{IMC}}

2. The CMB TF measurement is taken with the IMC locked and the slow LSC loop enabled (all 3 loop on). The measured TF will include IMC loop supression and suppression from the LSC loop, which is further suppressed by the IMC loop. The expression below is what you obtain when calculating i1/i2 in Attachment 1 (see derivation in Attachment 2):

G_{CMB}^{IMC_{on}, LSC_{on}} = \frac{G_{CMB}}{(1-G_{IMC})(1 - \frac{G_{LSC}}{1-G_{IMC}})} = \frac{G_{IMC}\,C_{CARM}\,F_{CMB}}{(1-G_{IMC})(1-G_{LSC}^{IMC_{on},CMB_{off}})}

Note that the final expression is in terms of known or measured quantitites. 

Attachment 1: IMG_5845.JPG
Attachment 2: IMG_5847.JPG
  17919   Wed Oct 25 09:23:31 2023 JCUpdateLSCC1SUS2 Crash Again


I came in this morning and began to fiddle with the Spectrogram. There was an error with the previous code and it doesn't have the ability to open anymore. After sometime of trying to fix the issue of the code, I ooked up and realized that IMC was not aligned. I checked WFS and the screen was frozen. After 10 min or so of searching, I realized the cds screen I was looking at was one from yesterday (it was frozen). After opening a new CDS screen, I noticed c1sus2 crashed again.


When using this Fencing technique on c1sus2, I haven't been able to avoid tohe issue of DK triggering on c1x04, and c1x01. Although, when I use the technique on C1LSC and C1ISCEX, it works perfectly fine. I may be doing this incorrectly though. Her are the steps I took this morning. Please let me know if I did something incorrectly.

  1. SSH into fb1 and run the following command " ./dolphin_ix_port_control.sh --disable 6 "
  2. ssh into c1sus and run "sudo reboot" This will return you back to fb1.
  3. ping c1sus2 until the machine stops responding.
  4. Once the machine stops responding, enable the dolphin port by running " ./dolphin_ix_port_control.sh --enable 6 "    This is done before the machine boots back up.

After I did this, DK trigger on onc1x04, and c1x01. I repeated the steps for both C1LSC and C1ISCEX and burtrestored the machines to yesterday at 17:19. After this, our system was brought back up to its original state before the c1sus2 crash.


  17918   Tue Oct 24 18:19:44 2023 yutaUpdateSUSActuator calibrations of BHD optics

[Begüm, Vittoria, Yuta]

We calibrated actuators for BHD optics under PRY and ITMY-LO fringe configurations using ITMY as a reference.

 - For calibration PR2 and PR3, we locked PRY using REFL55_I as an error signal and PRM as an actuator. We measured transfer functions from C1:LSC-ITMY_IN2 or C1:SUS-(PR2|PR3)_LSC_IN2 to C1:LSC-REFL11_I_ERR. Took the ratio of TFs with respect to ITMY (measured in 40m/17886) to calibrate PR2 and PR3 (Attachment #1 and #2).
 - For calibration of LO1, LO2, AS1, AS4, SR2, we locked ITMY single bounce vs LO fringe using BH55_Q as an error signal and BS as an actuator. We measured transfer functions from C1:LSC-ITMY_IN2 or C1:SUS-(XXX)_LSC_IN2 to C1:LSC-BH55_Q_ERR. Took the ratio of TFs with respect to ITMY to calibrate them (Attachment #3 and #4).

Summary of results:
 - Summarized in the floowing table. Results are in the unit of /f^2 nm/counts. If you divide the raw measurement (labeled Meas. nm/count) by "V2A" filters used for gain adjustments and are compensated by AoI or BS to MICH (factor of sqrt(2)) or PR2/3 to PRY (factor of 2), they are all in the range of 18-30 /f^2 nm/counts. LO2 have small actuation efficiency which was known because of too much pitch alignment offset. LO1, LO2, AS1, AS4 were calibrated before in 40m/17285, and the new measurements are consistent. PR2, PR3, SR2 calibration were done for the first time.

Susp. | Meas. nm/count | Gain adj. | AoI deg | nm/count
BS   +69.54 +2.325 45 +21.15
ITMX +14.73 +0.653  0 +22.55
ITMY +14.50 +0.659  0 +21.99
ETMX +12.20 +0.414  0 +29.47
ETMY +10.66 +0.480  0 +22.21
MC2  +2.27 +0.105  0 +21.62
PRM  -19.00 +0.773  0 -24.58
PR2  -36.19 +1.000  0 -18.09
PR3  -29.57 +1.000 45 -20.91
LO1  +28.25 +1.000  0 +28.25
LO2  +11.19 +1.000 45 +15.83
AS1  -25.92 +1.000  0 -25.92
AS4  +25.82 +1.000  0 +25.82
SR2  -16.63 +1.000 45 -23.52

 - Use these actuation efficiencies to dither each optic and see beam spot positions on them

Attachment 1: PR23ActuatorTF_20231024.pdf
Attachment 2: PR23ActuatorRatio_20231024.pdf
Attachment 3: BHDActuatorTF_20231024.pdf
Attachment 4: BHDActuatorRatio_20231024.pdf
  17917   Tue Oct 24 18:07:06 2023 yutaUpdateCDSc1sus2 crashed again

While locking ITMY single bounce vs LO fringe, c1sus2 crashed again (at around 17:20 local).cryingcryingcrying
c1sus2 rebooted with Dolphin thing, but it crashed c1lsc, c1sus, c1iscex, so we also had to reboot these machines.
Burt-restored all by /opt/rtcds/caltech/c1/Git/40m/scripts/cds/burtRestoreAndResetSUS.sh /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2023/Oct/24/16:19

ELOG V3.1.3-